OpenStudy (anonymous):

what is 258 over 301 i need to simplify it?

OpenStudy (anonymous):

258 over 301

OpenStudy (cwrw238):

write as prime factors: 258 = 2 * 129 129 = 3* 43 so 258 = 2*3*43

OpenStudy (cwrw238):

301 = 7 * 43

OpenStudy (anonymous):

what is the answer

OpenStudy (cwrw238):

so we get